# On Limiting Probability Distributions of Higher Order Markov Chains

### v1

@@ 2 Tensor Formulation and Matricization
Let us take a second order, i.e. $m=3$, chain $X$ on state space $\{1,2\}$ whose transition tensor $\cal P$ is as follows:
@@
For a second order chain, the following technical lemma can be seen in [25]. We now extend it to the general higher order chain $X$.
@@
Proof: Clearly, (2.6) holds when $u\geq 0$ and $v=0$. Assume now $u=0$ but $v\geq 1$. Note that on row $i_{1}$, $P^{(0)}$ has ones at, in multiindex form, columns $i_{1}j_{2}\ldots j_{m-1}$ for any $j_{2},\ldots,j_{m-1}\in S$ and zeros elsewhere. Hence,
$\begin{array}[]{rcl}(P^{(0)}Q^{v})_{i_{1},i_{2}\ldots i_{m}}&=&\displaystyle\sum_{j_{1}\ldots j_{m-1}}p^{(0)}_{i_{1},j_{1}\ldots j_{m-1}}q^{(v)}_{j_{1}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] &=&\displaystyle\sum_{j_{2}\ldots j_{m-1}}q^{(v)}_{i_{1}j_{2}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] &=&\displaystyle\sum_{j_{2}\ldots j_{m-1}}\Pr(X_{t+v}=i_{1},X_{t+v-1}=j_{2},\ldots,X_{t+v-m+2}=j_{m-1}\\ &&\hskip 60.00009pt|X_{t}=i_{2},X_{t-1}=i_{3},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] &=&p^{(v)}_{i_{1},i_{2}\ldots i_{m}},\end{array}$
@@
$\begin{array}[]{l}(P^{(u)}Q^{v})_{i_{1},i_{2}\ldots i_{m}}\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}p^{(u)}_{i_{1},j_{1}\ldots j_{m-1}}q^{(v)}_{j_{1}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}\Pr(X_{t+u+v}=i_{1}|X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1})\\ \hskip 40.00006pt\cdot\Pr(X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}\Pr(X_{t+u+v}=i_{1}|X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1},\\[8.00003pt] \hskip 120.00018ptX_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] \hskip 40.00006pt\cdot\Pr(X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}\Pr(X_{t+u+v}=i_{1},X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}\\[8.00003pt] \hskip 70.0001pt|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[10.00002pt] =p^{(u+v)}_{i_{1}i_{2}\ldots i_{m}},\end{array}$
@@
3 Limiting Probability Distribution
After setting the stage in the preceding section, we are ready to review the existing lines of inquiry concerning the limiting probability distribution of the $(m-1)$th order Markov chain $X$ on state space $S=\{1,2,\ldots,n\}$ with transition tensor $\cal P$.
The first approach assumes that on the reduced first order chain $Y$, if $\lim_{t\rightarrow\infty}y_{t}$ exists, then such a limit can be approximated by a rank-one symmetric tensor
$\underbrace{y\circ\cdots\circ y}_{m-1},$ (3.1)
where $y\in\mathbb{R}^{n}$ and $\circ$ denotes the outer product. This approach can be seen in most existing works on higher order chains; see, for example, [1, 2, 3, 4, 5, 6, 8, 12, 13, 20, 21]. It also leads to the interesting $Z$-eigenvector problem. It is known [21] that such $y$ can be determined as some normalized positive $Z$-eigenvector when $\cal P$ is irreducible. As pointed out in [26], however, the approximation in (3.1) may not be able to capture the true nature of the probability distribution $y_{t}$ and its limit.
The second approach [26] involves a two-phase power iteration scheme based on the connection between the higher order chain $X$ and its reduced first order chain $Y$. It assumes that $Q$ is regular, namely $\lambda=1$ is the only dominant eigenvalue of $Q$ on the unit circle. Then, $\lim_{t\rightarrow\infty}y_{t}$ turns out to be the eigenvector of $Q$ associated with $\lambda$ which is normalized with $\|\cdot\|_{1}$. Consequently, the limiting probability distribution of $X$ follows from (2.7). As we shall show, however, that the limiting probability distribution of $X$ may still exist despite $Q$ being non-regular.
The third approach [7] deals with the higher order chain $X$ by a certain approximation scheme to its transition probabilities. The development also hinges on a direct extension of the notion of recurrent states from the first order case to the higher order case. Such an extension has been shown [11] to be an issue, which turns out to be another difference between the higher order and the first order cases.
Motivated by the above studies we have reviewed, especially by [26], our goals in this paper are to further investigate the actual limiting probability distribution of the higher order chain $X$ as well as to establish a new sufficient condition that guarantees the existence of such a limiting probability distribution. This new condition uses the regularity of $\cal P$ instead of that of $Q$ and, therefore, strengthens the existing results in [26].
We now state our main result as follows.
@@ 3 Limiting Probability Distribution
Given an $(m-1)$th order Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S$, suppose that its transition tensor $\cal P$ is regular. Then, there exists $\pi\in\mathbb{R}^{n}$ satisfying $\pi_{i}>0$ for each $i$ and $\sum_{i=1}^{n}\pi_{i}=1$ such that $\lim_{k\rightarrow\infty}p^{(k)}_{ii_{2}\ldots i_{m}}=\pi_{i}$ for all $i_{2},\ldots,i_{m}\in S$, i.e.
$\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\circ\underbrace{e\circ\cdots\circ e}_{m-1}.$ (3.2)
@@
For the case of second order chains, a result similar to Theorem 3.1 can be found in [25].
Our next result confirms that $\pi$ in Theorem 3.1 is indeed the limiting probability distribution of the chain $X$ which contains the long-term probabilities for the chain to be in various states. Moreover, these probabilities are independent of the initial distribution of $X_{1},X_{2},\ldots,X_{m-1}$.
@@
Suppose that $X=\{X_{t}:t=1,2,\ldots\}$ is an $(m-1)$th order Markov chain on state space $S$ whose transition tensor is $\cal P$. If there exists a nonnegative $\pi\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}\pi_{i}=1$ and ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\circ\underbrace{e\circ\cdots\circ e}_{m-1}}$, then
@@
$\begin{split}\Pr(X_{m+k-1}=i)&=\sum_{i_{2},\ldots,i_{m}=1}^{n}\Pr(X_{m+k-1}=i,X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\[5.0pt] &=\sum_{i_{2},\ldots,i_{m}=1}^{n}\Pr(X_{m+k-1}=i|X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\ &\hskip 57.81621pt\cdot\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\[5.0pt] &=\sum_{i_{2},\ldots,i_{m}=1}^{n}p^{(k)}_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).\end{split}$
Pushing $k\rightarrow\infty$ and noting $\sum_{i_{2},\ldots,i_{m}=1}^{n}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m})=1$, the proof is complete. $\Box$
We comment here that Theorem 3.2 does not require $\pi$ to be positive. This will allow us to explore in future works limiting behavior of the chain $X$ in broader settings such as higher order absorbing chains.
From Theorems 3.1 and 3.2, we see that the limiting probability distribution of the chain $X$ can be investigated entirely within a tensor theoretic framework without resorting to the reduced first order chain or relying on the two-phase power iteration scheme as in [26].
Theorem 3.1 establishes a sufficient condition, that is, $\cal P$ being regular, for the existence of the limiting probability distribution of the chain $X$. In addition, it shows that in this case, the limiting probability distribution is positive. This extends the well-known result regarding first order chains as recalled in the introductory section.
@@
From the proof of Theorem 3.2, we can also see the role that the limit ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\circ\underbrace{e\circ\cdots\circ e}_{m-1}}$ plays so as to guarantee that the limiting probability distribution is independent of the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$. This result can be extended slightly as follows.
@@
Suppose that $X=\{X_{t}:t=1,2,\ldots\}$ is an $(m-1)$th order Markov chain on state space $S$ and with transition tensor $\cal P$. If there exists some tensor $\cal R$ of the same size as $\cal P$ such that $\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\cal R$, then for each $i\in S$, $\lim_{t\rightarrow\infty}\Pr(X_{t}=i)$ exists but, in general, it depends on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$.
@@
$\Pr(X_{m+k-1}=i)=\sum_{i_{2},\ldots,i_{m}=1}^{n}p^{(k)}_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).$
@@
$\lim_{t\rightarrow\infty}\Pr(X_{t}=i)=\sum_{i_{2},\ldots,i_{m}=1}^{n}r_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).$ (3.5)
@@
In view of Theorem 3.3, we now revisit Example 2.1.
@@
Our final result connects the limiting probability distribution with the dominant eigenvectors of $Q$. A similar result for a regular $Q$ is provided in [26] too.
@@
Let $\pi$ be the limiting probability distribution of an $(m-1)$th order regular Markov chain $X$ on state space $S$. Let $Q$ be the transition matrix of its reduced first order chain $Y$ and assume that $z\in\mathbb{R}^{N}$ is a normalized nonnegative right eigenvector of $Q$ associated with dominant eigenvalue $1$, i.e. $Qz=z$, $z\geq 0$, and $e^{T}z=1$. Then,
$\pi=P^{(0)}z,$ (3.6)
@@
Proof: According to (2.7), we have
$P^{(k)}z=P^{(0)}Q^{k}z=P^{(0)}z.$
@@
We remark that the above result holds even when $Q$ is non-regular and has multiple dominant eigenvectors associated with eigenvalues of modulus $1$. In fact, this may happen even when $\cal P$ is regular, as illustrated by the example below from [7].
@@
Take a second order chain on the state space $\{1,2,3,4\}$ whose transition tensor $\cal P$ has the form
@@
$\pi=P^{(0)}z^{(1)}=P^{(0)}z^{(2)}=\frac{1}{7}[2~{}2~{}2~{}1]^{T}.$

### v2

@@ 3 Main Results
Take a second order chain $X$ on state space $\{1,2\}$ whose transition tensor $\cal P$ is as follows:
@@
In the rest of this paper, we assume $m\geq 3$ unless otherwise stated.
For a second order chain, the following technical lemma can be seen in [25]. We now extend it to the more general higher order chain $X$. It should be pointed out, however, that such an extension from the second order case to the general higher order case is usually not straightforward, as evidenced by a shortage of results for the latter. A key reason for this situation lies in the fact that the general case is often far less tractable.
@@
Proof: Clearly, (3.1) holds when $u\geq 0$ and $v=0$. Assume now $u=0$ but $v\geq 1$. Note that on row $i_{1}$, $P^{(0)}$ has ones at, in multi-index form, columns $i_{1}j_{2}\ldots j_{m-1}$ for any $j_{2},\ldots,j_{m-1}\in S$ and zeros elsewhere. Hence,
$\begin{array}[]{rcl}(P^{(0)}Q^{v})_{i_{1},i_{2}\ldots i_{m}}&=&\displaystyle\sum_{j_{1}\ldots j_{m-1}\in T}p^{(0)}_{i_{1},j_{1}\ldots j_{m-1}}q^{(v)}_{j_{1}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] &=&\displaystyle\sum_{j_{2},\ldots,j_{m-1}\in S}q^{(v)}_{i_{1}j_{2}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] &=&\displaystyle\sum_{j_{2},\ldots,j_{m-1}\in S}\Pr(X_{t+v}=i_{1},X_{t+v-1}=j_{2},\ldots,X_{t+v-m+2}=j_{m-1}\\ &&\hskip 60.00009pt|X_{t}=i_{2},X_{t-1}=i_{3},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] &=&p^{(v)}_{i_{1},i_{2}\ldots i_{m}},\end{array}$
@@
$\begin{array}[]{l}(P^{(u)}Q^{v})_{i_{1},i_{2}\ldots i_{m}}\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}\in T}p^{(u)}_{i_{1},j_{1}\ldots j_{m-1}}q^{(v)}_{j_{1}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}\in T}\Pr(X_{t+u+v}=i_{1}|X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1})\\ \hskip 40.00006pt\cdot\Pr(X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}\in T}\Pr(X_{t+u+v}=i_{1}|X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1},\\[8.00003pt] \hskip 120.00018ptX_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] \hskip 40.00006pt\cdot\Pr(X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}\in T}\Pr(X_{t+u+v}=i_{1},X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}\\[8.00003pt] \hskip 70.0001pt|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[10.00002pt] =p^{(u+v)}_{i_{1}i_{2}\ldots i_{m}},\end{array}$
@@
The above lemma shows a connection between the power of $Q$ and the power of $\cal P$ in matricized form. It has several applications, including in the proof of Theorem 3.6 later. Here, we first prove a sufficient condition for $\cal P$ to be regular, which generalizes a similar result in [25] for the second order case:
@@
Let $\cal P$ be the transition tensor of the higher order chain $X$ and $Q$ be the transition matrix of its reduced first order chain $Y$. If $Q$ is regular, then so is $\cal P$.
Proof: Since $Q$ is regular, there exists $k\geq 1$ such that $Q^{k}>0$. The conclusion follows from (3.2). $\Box$
Theorem 3.1 provides a way of determining the regularity of $\cal P$ through $Q$, which allows us to make use of the available results in this regard [12]. In general, the converse of Theorem 3.1 is not true. We, however, have the following for the case ${\cal P}>0$:
THEOREM 3.2
For the $(m-1)$th order Markov chain $X$ on state space $S$ with transition tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]>0$, denote $Q^{k}=[q^{(k)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}]$ using multi-index form. Then
$q^{(m-1)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}=\underbrace{p_{i_{1}i_{2}\ldots i_{m-1}j_{1}}p_{i_{2}i_{3}\ldots i_{m-1}j_{1}j_{2}}\cdots p_{i_{m-1}j_{1}j_{2}\ldots j_{m-1}}}_{m-1}>0,$
which implies that $Q$ is regular.
Proof: We claim that for any $12\leq\ell\leq m-$ 1,
$\begin{array}[]{l}q^{(\ell)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}\\ =\left\{\begin{array}[]{cl}\underbrace{p_{i_{1}i_{2}\ldots i_{m-1}j_{m-\ell}}p_{i_{2}i_{3}\ldots i_{m-1}j_{m-\ell}j_{m-\ell+1}}\cdots p_{i_{\ell}i_{\ell+1}\ldots i_{m-1}j_{m-\ell}j_{m-\ell+1}\ldots j_{m-1}}}_{\ell},&\begin{array}[]{l}j_{1}=i_{\ell+1},\\ j_{2}=i_{\ell+2},\ldots,\\ j_{m-\ell-1}=i_{m-1};\end{array}\\ 0,&~{}{\rm otherwise}.\end{array}\right.\end{array}$ (3.3)
Note that the number of conditions for having a positive $q^{(\ell)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}$ decreases as $\ell$ increases. Any such condition is interpreted here as null when any subscript is out of range. Let us verify the above by induction.
When $\ell=2$,
$\begin{array}[]{rcl}q^{(2)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}&=&\displaystyle\sum_{k_{1}k_{2}\ldots k_{m-1}\in T}q_{i_{1}i_{2}\ldots i_{m-1},k_{1}k_{2}\ldots k_{m-1}}q_{k_{1}k_{2}\ldots k_{m-1},j_{1}j_{2}\ldots j_{m-1}}\\ &=&q_{i_{1}i_{2}\ldots i_{m-1},i_{2}i_{3}\ldots i_{m-1}j_{m-2}}q_{i_{2}i_{3}\ldots i_{m-1}j_{m-2},j_{1}j_{2}\ldots j_{m-1}}\end{array}$
since $k_{1}=i_{2},k_{2}=i_{3},\ldots,k_{m-2}=i_{m-1}$, and $k_{m-1}=j_{m-2}$; otherwise, either $q_{i_{1}i_{2}\ldots i_{m-1},k_{1}k_{2}\ldots k_{m-1}}=0$ or $q_{k_{1}k_{2}\ldots k_{m-1},j_{1}j_{2}\ldots j_{m-1}}=0$. Thus,
$q^{(2)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}=\left\{\begin{array}[]{cl}p_{i_{1}i_{2}\ldots i_{m-1}j_{m-2}}p_{i_{2}i_{3}\ldots i_{m-1}j_{m-2}j_{m-1}},&j_{1}=i_{3},j_{2}=i_{4},\ldots,j_{m-3}=i_{m-1};\\ 0,&{\rm otherwise}.\end{array}\right.$
Suppose now that the claim holds for any $22\leq\ell\leq m-$ 2. Then,
$\begin{array}[]{rcl}q^{(\ell+1)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}&=&\displaystyle\sum_{k_{1}k_{2}\ldots k_{m-1}\in T}q_{i_{1}i_{2}\ldots i_{m-1},k_{1}k_{2}\ldots k_{m-1}}q^{(\ell)}_{k_{1}k_{2}\ldots k_{m-1},j_{1}j_{2}\ldots j_{m-1}}\\ &=&p_{i_{1}i_{2}\ldots i_{m-1}j_{m-\ell-1}}q^{(\ell)}_{i_{2}i_{3}\ldots i_{m-1}j_{m-\ell-1},j_{1}j_{2}\ldots j_{m-1}},\end{array}$
where we note that $k_{1}=i_{2},k_{2}=i_{3},\ldots,k_{\ell+1}=i_{\ell+2},\ldots,k_{m-2}=i_{m-1}$, and, using the inductive hypothesis, $k_{m-1}=j_{m-\ell-1}$. Furthermore, using the inductive hypothesis again,
$\begin{array}[]{l}q^{(\ell+1)}_{i_{1}i_{2}\ldots i_{m-1},j_{1}j_{2}\ldots j_{m-1}}\\ =\left\{\begin{array}[]{cl}p_{i_{1}i_{2}\ldots i_{m-1}j_{m-\ell-1}}\underbrace{p_{k_{1}k_{2}\ldots k_{m-1}j_{m-\ell}}\cdots p_{k_{\ell}k_{\ell+1}\ldots k_{m-1}j_{m-\ell}j_{m-\ell+1}\ldots j_{m-1}}}_{\ell},&\begin{array}[]{l}j_{1}=k_{\ell+1},\\ j_{2}=k_{\ell+2},\ldots,\\ j_{m-\ell-1}=k_{m-1};\end{array}\\ 0,&~{}{\rm otherwise},\end{array}\right.\\ =\left\{\begin{array}[]{cl}\underbrace{p_{i_{1}i_{2}\ldots i_{m-1}j_{m-\ell-1}}p_{i_{2}i_{3}\ldots i_{m-1}j_{m-\ell-1}j_{m-\ell}}\cdots p_{i_{\ell+1}i_{\ell+2}\ldots i_{m-1}j_{m-\ell-1}j_{m-\ell}\ldots j_{m-1}}}_{\ell+1},&\begin{array}[]{l}j_{1}=i_{\ell+2},\\ j_{2}=i_{\ell+3},\ldots,\\ j_{m-\ell-2}=i_{m-1};\end{array}\\ 0,&~{}{\rm otherwise}.\end{array}\right.\par\end{array}$
The proof, therefore, is done. $\Box$
In what follows, $\otimes$ stands for the outer product. We now state our main result as follows.
THEOREM 3.3
Consider the $(m-1)$th order Markov chain on state space $S$. Suppose that its transition tensor $\cal P$ is regular. Then, there exists $\pi\in\mathbb{R}^{n}$ satisfying $\pi_{i}>0$ for each $i$ and $\sum_{i=1}^{n}\pi_{i}=1$ such that $\lim_{k\rightarrow\infty}p^{(k)}_{ii_{2}\ldots i_{m}}=\pi_{i}$ for all $i_{2},\ldots,i_{m}\in S$, i.e.
$\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\otimes\underbrace{e\otimes\cdots\otimes e}_{m-1}.$ (3.4)
@@
We comment that for the second order case, a result similar to Theorem 3.3 can be found in [25]. Our next result confirms that $\pi$ as in Theorem 3.3 is indeed the limiting probability distribution of the chain, which contains the long-term probabilities for the chain to be in various states. Moreover, these probabilities are independent of the initial probability distributions.
@@
Let $X=\{X_{t}:t=1,2,\ldots\}$ be an $(m-1)$th order Markov chain on state space $S$ whose transition tensor is $\cal P$. If there exists a nonnegative $\pi\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}\pi_{i}=1$ and ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\otimes\underbrace{e\otimes\cdots\otimes e}_{m-1}}$, then
@@
$\begin{split}\Pr(X_{m+k-1}=i)&=\sum_{i_{2},\ldots,i_{m}\in S}\Pr(X_{m+k-1}=i,X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\[5.0pt] &=\sum_{i_{2},\ldots,i_{m}\in S}\Pr(X_{m+k-1}=i|X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\ &\hskip 57.81621pt\cdot\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\[5.0pt] &=\sum_{i_{2},\ldots,i_{m}\in S}p^{(k)}_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).\end{split}$
Pushing $k\rightarrow\infty$ and noting $\sum_{i_{2},\ldots,i_{m}\in S}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m})=1$, the proof is now complete. $\Box$
We remark that Theorem 3.4 does not require $\pi$ to be strictly positive. This will allow us to explore in future works limiting behavior of the chain $X$ in other settings such as being a higher order absorbing chain.
We also point out that if there exists $y\in\mathbb{R}^{N}$ such that $\lim_{k\rightarrow\infty}Q^{k}=ye^{T}$, then, by Lemma 3.1, the mode-$1$ matricization of ${\cal P}^{(k)}$ satisfies $\lim_{k\rightarrow\infty}P^{(k)}=(P^{(0)}y)e^{T}$. This translates exactly to ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\otimes\underbrace{e\otimes\cdots\otimes e}_{m-1}}$, where $\pi=P^{(0)}y$ is the limiting probability distribution according to Theorem 3.4. We shall further connect the limiting probability distribution to an eigenvector of $Q$ later in this section.
Along with Theorem 3.4, Theorem 3.3 establishes a new sufficient condition, namely, $\cal P$ being regular, for the existence of the limiting probability distribution of the chain $X$. Moreover, it shows that in this case, the limiting probability distribution is positive. These extend the well-known results regarding first order chains as discussed in our introductory section.
@@
The proof of Theorem 3.4 demonstrates how ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\otimes\underbrace{e\otimes\cdots\otimes e}_{m-1}}$ can ensure that the limiting probability distribution does not hinge on the initial probability distributions. The result below deals with the case when $\lim_{k\rightarrow\infty}{\cal P}^{(k)}$ no longer has the special rank-one structure as in Theorem 3.4.
@@
Consider an $(m-1)$th order Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ with state space $S$ and transition tensor $\cal P$. If there exists a tensor $\cal R$ such that $\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\cal R$, then for each $i\in S$, $\lim_{t\rightarrow\infty}\Pr(X_{t}=i)$ exists. In general, however, such a limit depends on the initial probability distributions of $X$.
@@
$\Pr(X_{m+k-1}=i)=\sum_{i_{2},\ldots,i_{m}\in S}p^{(k)}_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).$
@@
$\lim_{t\rightarrow\infty}\Pr(X_{t}=i)=\sum_{i_{2},\ldots,i_{m}\in S}r_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).$ (3.7)
@@
Now, in view of Theorem 3.5, we revisit Example 3.1 to illustrate how $\lim_{t\rightarrow\infty}\Pr(X_{t}=i)$ may depend on the initial probability distributions.
@@
Our final result connects the limiting probability distribution with the dominant eigenvectors of $Q$. A similar result can be found in [26].
@@
Suppose that $\pi$ is the limiting probability distribution of an $(m-1)$th order regular Markov chain $X$ on state space $S$. Let $Q$ be the transition matrix of its reduced first order chain. Suppose that $y\in\mathbb{R}^{N}$ is a normalized nonnegative right eigenvector of $Q$ associated with the dominant eigenvalue $\lambda=1$, i.e., $Qy=y$, $y\geq 0$, and $e^{T}y=1$, then,
$\pi=P^{(0)}y,$ (3.8)
@@
Proof: According to (3.2) in Lemma 3.1, we have
$P^{(k)}y=P^{(0)}Q^{k}y=P^{(0)}y.$
@@
We comment that, unlike in [26], the above result holds even when $Q$ is non-regular and has multiple dominant eigenvectors associated with eigenvalues of modulus $1$. In fact, this may happen even when $\cal P$ is regular, as illustrated by the example below from [7].
@@
Consider a second order chain on state space $\{1,2,3,4\}$, whose transition tensor $\cal P$ has the form
@@
$\pi=P^{(0)}y^{(1)}=P^{(0)}y^{(2)}=\frac{1}{7}[2~{}2~{}2~{}1]^{T}.$
4 Conclusion
In the paper, motivated by the relevant existing works, we have developed several useful properties regarding the limiting probability distribution of a general higher order Markov chain. In particular, we have shed light on the key role the regularity of the transition tensor plays. We have also provided a way of determining the regularity via the transition matrix of the reduced first order chain. Moreover, we have connected the limiting probability distribution to some eigenvector arising on the reduced first order chain, even when this chain is non-regular.